ROTATIONAL SPECTRUM AND LARGE AMPLITUDE MOTIONS OF 3,4-, 2,5- and 3,5-DIMETHYL-BENZALDEHYDE

\underline{I. KLEINER}, Laboratoire Interuniversitaire des Systèmes Atmosphériques, CNRS et Universités Paris Diderot et Paris Est, 61 av. Général de Gaulle, 94010, Créteil, France; M. TUDORIE, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, 50 av. F-D. Roosevelt, 1050 Bruxelles, Belgique; M. JAHN, J.-U. GRABOW, Gottfried-Wilhelm-Leibniz-Universität, Institut für Physikalische Chemie und Elektrochemie, Lehrgebiet A, Callinstraße 3A, D-30167 Hannover, Germany; M. GOUBET, Laboratoire PhLAM, UMR 8523 CNRS, Bât. P5, Université des Sciences et Technologies de Lille 1, F-59655 Villeneuve d’Ascq, France.

The microwave spectra of the 3,4-, 2,5- and 3,5-Dimethyl-Benzaldehyde (DMBA) molecules have been recorded for the first time in the 2-26.5 GHz frequency range, using the COBRA-FTMW spectrometer in Hannover, with an instrumental uncertainty of 0.5 kHz for unblended lines. The experimental assignments and fits are supplemented by ab initio quantum chemical calculations,conformational energy landscape, and dipole moment components. The analysis of the spectra for the three isomers are in progress. The latest results, including spectroscopic constants and large amplitude motion parameters, will be presented. This investigation^{a} follows the study of the spectra of the 4-Methyl-Benzaldehyde molecule^{b}. The DMBA isomers belong to a similar series of molecules formally obtained by adding a second methyl group at the aromatic ring. These molecules serve as prototype systems for the development of the theoretical model of asymmetric top molecules having C_{s} symmetry while containing two inequivalent methyl tops (C_{3v}), exhibiting different barrier heights and coupling terms to methyl internal rotation. Thus, the DMBA isomers represent benchmark species for testing the two-top internal rotors BELGI program written recently^{c}

aSupported by the ANR-08-BLAN-0054 contract (France), the Deutsche Forschungsgemeinschaft, and the Land Niedersachsen (Germany).

bH. Saal, W. Caminati, I. Kleiner, A. R. Hight-Walker, J. T. Hougen, J.-U. Grabow, to be published.

cM. Tudorie, I. Kleiner, J. T. Hougen, S. Melandri, L. W. Sutikdja, W. Stahl, J. Mol. Spectrosc., 269 (2011), 211-225